### v1

# Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning

Abstract
Few-shot named entity recognition can identify new types of named entities based on a few labeled examples. Previous methods employing token-level or span-level metric learning suffer from the computational burden and a large number of negative sample spans. In this paper, we propose the Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning (MsFNER), which splits the general NER into two stages: entity-span detection and entity classification. There are 3 processes for introducing MsFNER: training, finetuning, and inference. In the training process, we train and get the best entity-span detection model and the entity classification model separately on the source domain using meta-learning, where we create a contrastive learning module to enhance entity representations for entity classification. During finetuning, we finetune the both models on the support dataset of target domain. In the inference process, for the unlabeled data, we first detect the entity-spans, then the entity-spans are jointly determined by the entity classification model and the KNN. We conduct experiments on the open FewNERD dataset and the results demonstrate the advance of MsFNER.
Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning
Peipei Liu1,2, Gaosheng Wang1,2, Ying Tong3, Jian Liang3, Zhenquan Ding1thanks: The corresponding author, Hongsong Zhu1,2 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences 3Jiangsu Provincial Department of Public Security, China {liupeipei, wanggaosheng, dingzhenquan, zhuhongsong}@iie.ac.cn
1 Introduction
Named Entity Recognition (NER) is a fundamental task in Natural Language Processing (NLP), involving the identification and categorization of text spans into predefined classes such as people, organizations, and locations Yadav and Bethard (2018); Li et al. (2022). Although traditional deep neural network architectures have achieved success in the fully supervised named entity recognition (NER) task with sufficient training data Lample et al. (2016); Ma and Hovy (2016), they are difficult to adapt to the dynamic nature of real-world applications, which require the model to quickly adjust itself to new data or environmental changes. In this case, researchers have proposed the few-shot named entity recognition (Few-shot NER) to explore entity recognition with limited labeled data Fritzler et al. (2019). Few-shot NER enables existing models to quickly transfer learned knowledge and adapt to new domains or entity classes.
Specifically, the Few-shot NER model is first trained on the source domain dataset $D_{source}=(S_{source},Q_{source})$ and then the trained model is transferred to new target domain dataset $D_{target}=(S_{target},Q_{target})$ to infer for $Q_{target}$ Snell et al. (2017). $S_{target}$ represents the support dataset comprising $N$ entity types ($N$-way), and each type is exemplified by $K$ annotated examples ($K$-shot). The $Q_{target}$ is the query dataset with the same entity types as the support dataset. Few-shot paradigm can offer a flexible and cost-effective solution to the adaptability challenge, making it a focal point of research to enhance the performance of NER systems in scenarios with limited labeled data or emerging entity types.
Currently, there are two mainstream researches for Few-shot NER: token-level Fritzler et al. (2019); Hou et al. (2020); Yang and Katiyar (2020); Das et al. (2022) and span-level metric learning methods Wang et al. (2022a); Yu et al. (2021); Ma et al. (2022). In token-level methods, each token is assigned an entity label based on its distances from the prototypes of entity classes or the support tokens. However, these approaches often have high computational costs and fail to maintain the semantic integrity of entity tokens, leading to increased susceptibility to interference from non-entity markers. On the other hand, although span-level methods can mitigate the partial issues associated with token-level approaches by evaluating entities as spans, all possible spans are enumerated would result in an N-square complexity and an increase in noise from a large number of negative samples.
Considering the challenges, we hope to solve the following problems: 1) To improve the Few-shot NER identification efficiency, how can we encourage the semantic divergence between entities and non-entities to determine effective entity spans? 2) To improve the entity span classification, how can we control and coordinate the semantic distance of different entity types to make the semantic representations of entities within the same types be proximate while those in disparate types be distant?
In this paper, we propose a hybrid multi-stage decoding approach for few-shot NER with contrastive learning (i.e., MsFNER). Specifically, MsFNER splits the general NER into two stages and this paper describes them by three processes: training process, finetuning process, and inference process. The training process consists of two objectives: training the best entity span detection model and the best entity classification model separately based on the supervised support dataset of source domain. In this process, we regard the training task as a meta-learning task and train entity detection model by employing Model-Agnostic Meta-Learning(MAML)Finn et al. (2017), and train entity classification model through the entity-aware contrastive learning and MAML-ProtoNetMa et al. (2022). In the finetuning process, we finetune the trained entity detection model and entity classification model on the support dataset of target domain with the same pattern as training process. In the inference process, we deploy the KNN and two finetuned models on the query dataset of target domain to predict entity spans and their types, and there are four phases for prediction. In the first phase, we build a key-value datastore $\mathcal{D}_{knn}$ on the support dataset $S_{target}$, where key is the entity representation and value is the corresponding label. Subsequently, we apply the entity detection model to get entity spans. With the entity detection model, we could reduce the impact of negative samples and reduce computational complexity. After that, we pass the representations of detected entity spans into the entity classification model to produce their predicted type distribution $p_{soft}$ while we also feed the representations into KNN to obtain the predicted results $p_{knn}$ based on the $\mathcal{D}_{knn}$. In the final phase, we combine $p_{soft}$ and $p_{knn}$ to obtain final types for prediction entities.
We summarize our contributions as follows:
  • We propose a hybrid multi-stage decoding for few-shot NER with entity-aware contrastive learning, in which we first detect entity spans to improve efficiency and then employ contrastive learning or KNN to augment entity classification performance.
  • Experimental results show that our model achieves new SOTA performance compared with previous works. We also evaluate the few-shot NER task on LLM ChatGPT and the experimental results show that our model outperforms ChatGPT in terms of performance and efficiency.
2 Methodology
In this section, we will introduce the details of MsFNER. The following content is arranged according to the three stages of the model.
2.1 Training Process
In this process, we introduce the separate process of training and finding the optimal entity span detection model and entity classification model.
Figure 1: The schematic diagram of the training process.
2.1.1 Entity Span Detection (ESD)
[…]
For a given sentence $x$ with $n$ tokens $x$=$(x_{1},x_{2},...,x_{n})$ in $D_{source}$, we first encode it using the pre-trained language model BERT. After that, we can obtain the contextualized representations $h=(h_{1},h_{2},...,h_{n})$ for all tokens.
The sequence token representations are then sent into a CRF to detect entity spans. Finally, the training loss in $D_{source}$ can be depicted as:
$L_{ESD}=-\sum logP(y|x)$ (1)
$P(y|x)=\frac{\prod_{i=1}^{|x|}\varphi_{i}(y_{i-1},y_{i},x)}{\sum_{y^{{}^{\prime}}}\prod_{i=1}^{|x|}\varphi_{i}(y_{i-1}^{{}^{\prime}},y_{i}^{{}^{\prime}},x)}$ (2)
where $\varphi_{i}(y_{i-1},y_{i},x)$ and $\varphi_{i}(y_{i-1}^{{}^{\prime}},y_{i}^{{}^{\prime}},x)$ are potential functions.
We regard the training process as a meta-learning task and divide the $D_{source}$ into many individual meta learning tasks. Then, we train the best entity-span detection model by MAML.
2.1.2 Entity Classification
For an entity $e_{k}=(x_{f},...,x_{f+l})$, where $f$ is the first token index and $f+l$ is the last token index in the sequence, we employ the max-pooling to get its representation when $l>=1$ :
$\hat{e}_{k}=max(h_{f},...,h_{f+l})$ (3)
Contrastive learning could enhance the consistency between entities within the same types and widen the distance between entities belonging to different types. As $e_{k}$ is the sample with certain type from the supervised $D_{source}$, we consider using the supervised contrastive learning to deepen the distinctiveness of entity types and augment representations of entities for entity classification improvement.
[…]
$L_{CL}=\sum_{j}-\frac{1}{|P(j)|}\sum_{p\in P(j)}log\frac{exp(sim(z^{j},z^{p})/\tau)}{\sum_{a\in A(j)}exp(sim(z^{j},z^{a})/\tau)}$ (4)
where P(j) is the positive set whose entities are from the same type with $e_{j}$, $z^{j}$ is the result of transforming $\hat{e}_{j}$ through a projection network MLP, A(j) is the set containing all N entities except for $e_{j}$, sim(·,·) denotes the KL-divergence for $K$-shot ($K$>1) but the squared euclidean distance for $K$-shot ($K$=1) referring to Das et al. (2022), and $\tau$ is a temperature hyperparameter.
After pulling entities of the same type together and entities of different types farther apart in the semantic space with contrastive learning, we discard the projection network at classification time following the previous works of contrastive learning Chen et al. (2020); Khosla et al. (2020); Liu et al. (2023). As a consequence, we use the entity representation $\hat{e}_{k}$ to compute the prototype $c_{t}(\mathcal{S})$ of $t_{th}$ type Snell et al. (2017), and predict the type distribution of $e_{k}$ by:
$p_{soft}(e_{k})=\frac{\exp\left(-d\left(c_{t}(\mathcal{S}),\hat{e}_{k}\right)\right)}{\sum_{ii\in\phi}\exp\left(-d\left(c_{ii}(\mathcal{S}),\hat{e}_{k}\right)\right)}$ (5)
[…]
2.2 Finetuning Process
Model finetuning aims to enable the model to adapt to new knowledge types during the domain transfer process. We finetune the trained models for entity detection and entity classification on the support dataset of the target domain $S_{target}$, using the same method as training process.
2.3 Inference Process
Figure 2: The schematic diagram of the inference process for entity classification.
In the inference process, we test the finetuned model on query dataset $Q_{target}$, and we will present this stage by four parts:
(1) We compute the representation $\hat{e}_{k}$ of each true entity in the support dataset $S_{target}$ by Eq.3 and build a key-value datastore $\mathcal{D}_{knn}$ where key is the entity representation and value is entity type.
(2) We input the query sentence $x^{\prime}$ into the finetuned best model for entity span detection (ESD) to get all entities. Specifically, we apply the Viterbi algorithm Forney (1973) for decoding, and derive the entities from the maximum sequence probability of Eq.2:
$y^{*}=argmax_{y\in\mathbb{Y}}P(y|x^{\prime})$ (7)
where $\mathbb{Y}$ is a set of all possible label sequences.
Models FewNERD-INTRA FewNERD-INTER
1-2 shot 5-10 shot 1-2 shot 5-10 shot
5 way 10 way 5 way 10 way 5 way 10 way 5 way 10 way
ProtoBERT† 23.45±0.92 19.76±0.59 41.93±0.55 34.61±0.59 44.44±0.11 39.09±0.87 58.80±1.42 53.97±0.38
NNShot† 31.01±1.21 21.88±0.23 35.74±2.36 27.67±1.06 54.29±0.40 46.98±1.96 50.56±3.33 50.00±0.36
StructShot† 35.92±0.69 25.38±0.84 38.83±1.72 26.39±2.59 57.33±0.53 49.46±0.53 57.16±2.09 49.39±1.77
CONTAINER† 40.436 33.84 53.70 47.49 55.95 48.35 61.83 57.12
ESD† 41.44±1.16 32.29±1.10 50.68±0.94 42.92±0.75 66.46±0.49 59.95±0.69 74.14±0.80 67.91±1.41
MAML-ProtoNet† 52.04±0.44 43.50±0.59 63.23±0.45 56.84±0.14 68.77±0.24 63.26±0.40 71.62±0.16 68.32±0.10
ChatGPT‡ 61.86 55.61 64.17 54.79 61.05 57.94 64.34 58.58
MsFNER (Ours) 54.25±0.34 46.69±0.5 66.57±0.29 58.70±1.39 72.91±0.34 66.34±0.65 78.41±0.30 72.06±0.11
w/o cl* 65.72±0.33 57.33±0.08 77.84±0.42 71.23±0.47
w/o KNN 53.86±0.47 46.24±0.29 66.04±0.65 57.56±0.41 72.75±0.29 65.38±0.72 77.99±0.25 71.92±0.14
Table 1: F1 scores with standard deviations on FewNERD. †denotes the results reported in Ma et al. (2022). ‡are the results we produce by LLM. * means the abbreviation of ‘contrastive learning’. The best results are in bold.
(3) Firstly, we compute the representation $\hat{e}^{\prime}_{k}$ of each detected entity ${e}^{\prime}_{k}$ by Eq.3. Then, we input $\hat{e}^{\prime}_{k}$ into two independent modules: on one hand, we feed the $\hat{e}^{\prime}_{k}$ into KNN to obtain the predicted result $p_{knn}$ based on the $\mathcal{D}_{knn}$ Wang et al. (2022b), and on the other hand we pass the $\hat{e}^{\prime}_{k}$ into the finetuned best model for entity classification to produce its predicted type distribution $p_{soft}$ by Eq.5.
(4) The final prediction type of the detected entity ${e}^{\prime}_{k}$ is jointly determined by KNN and the entity classification model:
$p(y|e^{\prime}_{k})=\lambda p_{knn}(y|e^{\prime}_{k})+(1-\lambda)p_{soft}(y|e^{\prime}_{k})$ (8)
where the $\lambda$ is a hyper-parameter that makes a balance between KNN and softmax.
3 Experiments
3.1 Experiments Setup
The details can be found at A.1.
3.2 Main results
Table 1 illustrates the main results of different methods on FewNERD dataset. The results demonstrate that MsFNER significantly outperforms all the previous state-of-the-art approaches. MsFNER attains average F1 improvements of 2.65 and 4.44 on INTRA and INTER, respectively, compared to the previous best method, MAML-ProtoNet. In the 5-way 5-10 shot setting on the INTER dataset, we achieve the most substantial improvement over the MAML-ProtoNet method, reaching 7.79 points. Additionally, we perform the few-shot task employing ChatGPT, where MsFNER surpasses ChatGPT in F1 performance by an average score of 8.465 in the 5-10 shot setting. Furthermore, MsFNER surpasses ChatGPT by an average of 10.13 percentage points in the 1-2 shot setting in the INTER dataset. Overall, regardless of whether it is ChatGPT or small models, increasing the number of $K$-shot yields superior performance, while an increase in the number of $N$-way results in worse performance.
3.3 Ablation Study
To investigate the contributions of different modules of MsFNER, we conduct the ablation study by removing each of them at a time to observe the performance of our model. The results are shown at the bottom of Table 1. Firstly, we remove the contrastive learning module. In the 1 shot experiment of this setting, KNN is unnecessary as there is only one sample in the support dataset. We can see that the removal of contrastive learning results in a decrease of 0.905 average score, which indicates the significant impact and necessity of contrastive learning for enhancing representation. Secondly, we remove the KNN module and reserve the contrastive learning module. The removal of KNN drops 0.523 average score. We can see that the impact of contrastive learning and KNN increases with the number of $N$-way and $K$-shot. Removing either of them will lead to a significant decrease in the performance of MsFNER.
3.4 Model Efficiency
To evaluate the efficiency of MsFNER, we compute the average inference time of MsFNER and ChatGPT in different settings. More details can be found at A.3.
4 Conclusion
In this paper, we propose the MsFNER, aiming to improve the performance and efficiency of few-shot NER. The MsFNER can be achieved by three stages: training, finetuning and inferring. We evaluate MsFNER on the open FewNERD dataset, and results show the advance of MsFNER compared with previous few-shot NER methods and LLM.
5 Limitations
  • In practical, the effectiveness of Few-shot NER is seriously affected due to the data sparsity and data imbalance.
  • The abundant external knowledge is very helpful for few-shot NER, which is worth exploring and studying.
6 Ethical Considerations
All the data we get is the open source, and there is no theft. We guarantee the originality, research ethics, societal impact and reproducibility of our work.
Appendix A Appendix
A.1 Experiments Setup
We conduct experiments on the Tesla V100 GPU.
A.1.1 Datasets
We conduct experiments on the FewNERD dataset Ding et al. (2021) which is annotated with 8 coarse-grained and 66 fine-grained entity types. We evaluate our method on the two settings of FewNERD: 1) FewNERD-INTRA, which pertains to a scenario wherein entities within the training set (source domain), validation set, and test set (target domain) are exclusively associated with distinct coarse-grained types. 2) FewNERD-INTER, where only fine-grained entity types that do not intersect in different sets. Furthermore, we follow the settings of ESD Wang et al. (2022a) and adopt $N$-way $K\sim 2K$-shot sampling method to construct tasks.
A.1.2 Parameter Settings
We use grid search for hyperparameter settings and we train our model for 1,000 steps and choose the best model on validation dataset for testing. We use the pretrained language model uncased BERT-base as our encoder. In the entity span detection module, we adopt the BIOES tags. We use the AdamW optimizer with a learning rate of 3e-5 and 0.01 linear warmup steps. The $K$ in KNN is set to 10 and the max-loss coefficient $\lambda$ of KNN is 0.1 when adding KNN. The batch size is set to 32, the max sequence length is set to 128 and we use a dropout probability of 0.2.
Models FewNERD-INTRA FewNERD-INTER
1-2 shot 5-10 shot 1-2 shot 5-10 shot
5 way 10 way 5 way 10 way 5 way 10 way 5 way 10 way
MsFNER 0.3442 0.3808 0.5616 1.0546 0.3482 0.3806 0.6708 1.2266
ChatGPT 4.5581 5.4897 8.1842 9.3481 3.2633 8.0579 4.7755 5.3409
Table 2: The comparison of inference time between MsFNER and ChatGPT.
A.1.3 Baselines
We compare MsFNER with popular baselines as follows:
  • ProtoBERT Fritzler et al. (2019) employs the prediction of query labels by leveraging the similarity between the BERT hidden states derived from the support set and the query tokens.
  • CONTAINER Das et al. (2022) employs token-level contrastive learning to train BERT as the token embedding function. Subsequently, BERT undergoes finetuning on the support set, followed by the application of a nearest neighbor method during the inference phase.
  • NNShot Yang and Katiyar (2020) bears resemblance to ProtoBERT in its methodology, as it conducts predictions utilizing nearest neighbor algorithms.
  • StructShot Yang and Katiyar (2020) incorporates an augmented Viterbi decoder into the inference process atop NNShot’s framework.
  • ESD Wang et al. (2022a) formulates few-shot sequence labeling as a span-level matching problem and decomposes it into span-related procedures.
  • MAML-ProtoNet Ma et al. (2022) proposes a decomposed meta-learning approach for few-shot NER, which tackles few-shot span detection and entity typing using MAML to facilitate rapid adaptation to novel entity classes and proposes MAML-ProtoNet for improved representation of entity spans.
  • ChatGPT chatgpt (2023) is an advanced neural language model developed by OpenAI, which is trained on diverse corpora of text and could generate human-like text.
A.2 ChatGPT Prompt Template
In this section, we introduce the ChatGPT few-shot NER prompt template that we use. The prompt is composed of 3 parts: task description, few-shot cases, and input queries.
In the task description, we initially elucidate the few-shot Named Entity Recognition (NER) task required from ChatGPT. Subsequently, we provide ChatGPT with explicit instructions for implementing this task, including delineation of the entity types involved. To facilitate easier interpretation of model output, we utilize a structured output JSON format, a domain in which ChatGPT exhibits proficiency. The template is as follows:
Your task is to perform fewshot Named Entity Recognition. You could perform the task by the following actions:
  1. 1.
    Do named entity recognition task to recognize the entities. Entity type is defined as \$ent_list
  2. 2.
    Check if the entity type is in the definition list.
  3. 3.
    Output a JSON object that contains the following keys: text, entity_list. Note that each item in entity_list is a dictionary with "entity" and "type" as keys.
Here are the given few-shot cases: ‘‘‘\$case_input‘‘‘
Output: \$case_output
Recognize the entities in the text delimited by triple backticks.: ‘‘‘\$input_text‘‘‘
In the template above, the \$ent_list denotes the list of entity types corresponding to $N$-way. The \$case_input comprises the sentences in the support set corresponding to $K$-shot, while the \$case_output represents the labels in the support set. Since ChatGPT is better at structured output, we use JSON format for output, with each output format being: {"entity": entity text ; "type": entity type}. The output will be merged into a list. The \$input_text denotes the query data. When we input this prompt to ChatGPT, it will return a list of relevant entity-type dictionaries.
A.3 Model Efficiency
To evaluate the efficiency of MsFNER, we compute the average inference time of MsFNER and ChatGPT in different settings. And the results are shown in Table 2, where we can see that MsFNER is on the millisecond level while the ChatGPT is on the second level. The time consumption is related to the amount of input data. In addition, the time consumption of ChatGPT will increase due to network impact.
A.4 training process
As the Figure 3 shows, the training loss on training dataset is smoothly declining within the same source types. But when finetuning on the validation set, there are some fluctuations due to the cross domain transfer of knowledge. And with the training time increasing, the overfitting is occurred slightly on validation dataset because of few supervised samples.
Figure 3: The performance of training process on FewNERD 5-way 5~10-shot. (a) is the evaluation F1 score when finetuning on the validation dataset. (b) is the training loss on train dataset.